sustitucion
2a+4b=2
5a+b=-13

Answer by addingup(3677)

2a+4b=2 (1)
5a+b=-13 (2)
``````````````````````````````
Divide all terms in equation (1)by 2:
a+2b=1
Now subtract 2b from both sides:
a = -2b+1 I'll use this value for a next
``````````````````````````````
In equation (2), substitute for the value of a:
5(-2b+1)+b=-13
-10b+5+b = -13
-9b= -18 Divide both sides by -9, and remember that -/- = +
b = 2
Now find the value of a. Take any of the two original equations, I'll use (1):
2a+4(2)=2
2a+8=2
2a=-6
a= -3